# Mod $\ell$ non-vanishing of self-dual Hecke $L$-values over CM fields and applications

### v1

@@ 3.2. Test vector
For a self-dual pair $(\pi,\chi)$, we have $\dim{\mathrm{Hom}}_{K_{{\mathbb{A}}}^{\times}}(\pi,\chi^{-1})\leq 1$ and a criterion for the equality is given by (3.2). Following Gross and Prasad [18], $f\in\pi$ is said to be a test vector for ${\mathrm{Hom}}_{K_{{\mathbb{A}}}^{\times}}(\pi,\chi^{-1})$ if any basis of ${\mathrm{Hom}}_{K_{{\mathbb{A}}}^{\times}}(\pi,\chi^{-1})$ takes a non-zero value on $f$. This is essentially a local notion: $f$ is a test vector if it is so locally101010defined analogously for any place. The Waldspurger formula [58, 61] links the Rankin–Selberg $L$-value
@@
with $K^{\times}$-toric period of a test vector in $\pi$ against $\chi$.
This subsection describes an explicit uniform test vector for the pairs $(\pi,\chi)$ as $\chi$ varies in the family ${\mathcal{X}}_{S_{0}}$ of Hecke characters. The next subsection describes an explicit relation between the $L$-values $L(1/2,\pi_{K}\otimes\chi)$ and toric periods of the test vector.
@@ 3.2.1. Definitions
For a finite place $v$, recall that $M_{0}({\mathfrak{p}}_{v}^{n})_{v}$ is the Eichler of discriminant ${\mathfrak{p}}_{v}^{n}$. Put $U_{0}({\mathfrak{p}}_{v}^{n})_{v}=M_{0}({\mathfrak{p}}_{v}^{n})_{v}^{\times}$. Recall that we have fixed an embedding $\iota:K\hookrightarrow B$ and a family of embeddings $\iota_{\varsigma^{({\mathfrak{c}})}}:K_{{\mathbb{A}}}^{\times}\hookrightarrow G({\mathbb{A}})$ in §2.2.1. Let $S_{1}$ denote the set of places dividing the conductor $\mathfrak{N}$ and $S_{0}$ a finite set of places at which $B$ splits.
@@ 3.2.2. Existence
    If $v\notin S_{0}$, then $R_{v}\cap\iota_{\varsigma_{v}}K_{v}=\iota_{\varsigma_{v}}O_{K_{v}}$, and if $v\in S_{0}$ and $\mathfrak{N}_{v}|{\mathfrak{c}}_{v}={\mathfrak{p}}_{v}^{c_{v}}$, then
    $R_{v}\cap\iota_{\varsigma_{v}^{(c_{v})}}O_{K_{v}}=\iota_{\varsigma_{v}^{(c_{v})}}O_{K_{v},{\mathfrak{c}}_{v}}.$
@@
Part i) just follows from [9, Lem. 3.3].
@@
The first part follows from newform theory and the Tunnell–Saito theorem (cf. [9, Prop. 3.7]). For the second part, we only need to consider the case that $v|\mathfrak{N}^{-}$ but $v\notin S_{0}$, which is the content of [9, Prop. 3.8]. ∎
@@
For any $\chi\in{\mathcal{X}}_{S_{0}}$ with conductor $\mathfrak{c}$}, the form
$f^{\varsigma^{({\mathfrak{c}})}}\in{\mathrm{Hom}}_{K_{{\mathbb{A}}}^{\times}}(\pi,\chi^{-1})$
is a test vector.
@@
For $v\nmid\mathfrak{N}^{-}$ or $v\in S_{0}$, note that $B_{v}$ splits, $f_{v}$ is the new vector and $\chi_{v}$ is unramified. So $f_{v}$ is a test vector by the condition $(a)$ in Definition 3.5. For $v|\mathfrak{N}^{-}$ but $v\nmid S_{0}$, one has $\chi_{v}=1$ and the assertion follows from definition. Lastly, for $v|S_{0}$ but $v\nmid\mathfrak{N}^{-}$, the assertion follows from the fact that under the condition $(b)$ in Definition 3.5, the local newform is a test vector for $\chi_{v}$. ∎
@@
Note that if $S_{0}$ is coprime to $\mathfrak{N}^{-}$, then the test vector $f$ does not depend on $S_{0}$.
@@
For any place ${\mathfrak{r}}|D_{K/F}$ of $F$, suppose that $ord_{{\mathfrak{r}}}(\mathfrak{N}_{\mathfrak{r}})>{\mathrm{ord}}_{{\mathfrak{r}}}(D_{K_{\mathfrak{r}}/F_{\mathfrak{r}}})$. Then ${\rm{N}}(R_{v}^{\times})={\rm{N}}(O_{K_{v}}^{\times})$ for any finite place $v$.
@@ 3.3.1. Setting
Let $B$ be a quaternion algebra over a number field $F$ and $G=B^{\times}$. Let $\pi$ be an irreducible cuspidal automorphic representation of $G({\mathbb{A}})$ with trivial central character and $\sigma$ its base change to ${\mathrm{GL}}_{2}({\mathbb{A}})$. Let $K$ be a quadratic field extension of $F$ with an embedding $K\hookrightarrow B$ and $\chi$ a finite order Hecke character over $K$ such that
@@
with the Rankin–Selberg $L$-value $L(1/2,\pi_{K}\otimes\chi)$.
@@
For $\phi$ the ${\mathrm{GL}}_{2}(F)$-newform of level
@@
associated to $\pi$, the Petersson norm $(\phi,\phi)_{U_{0}(\mathfrak{N})}$ is defined by the inner product on $B^{\times}\backslash{\mathbb{H}}^{\Sigma}\times\widehat{B}^{\times}/(\widehat{F}^{\times}U_{0}(N))$ arising from the invariant measure $dxdy/y^{2}$ on each component $\mathbb{H}$}, where $\mathbb{H}$} denotes the upper half plane.
Recall that $\widehat{R}\cap\iota_{\varsigma^{({\mathfrak{c}})}}\widehat{O}_{K}=\iota_{\varsigma^{({\mathfrak{c}})}}\widehat{O}_{K,{\mathfrak{c}}}$. For a Hecke character $\chi\in{\mathcal{X}}_{S_{0}}$ of conductor $\mathfrak{c}$}, the toric period $P_{\chi}(\pi(\varsigma^{({\mathfrak{c}})})f)$ with respect to embedding $\iota$ is essentially given by
$P(\varsigma^{({\mathfrak{c}})},\varphi,\chi):=\sum_{[a]\in{\rm G}_{\mathfrak{c}}}\chi(a)\langle{x_{0},\varphi(x_{\mathfrak{c}}(a))}\rangle_{k-2\Sigma},\quad f=\Psi(x_{0}\otimes\varphi)$
where ${\rm G}_{{\mathfrak{c}}}:=K^{\times}\widehat{F}^{\times}\backslash\widehat{K}^{\times}/\widehat{O}_{K,{\mathfrak{c}}}$ (cf. [9, Lem. 2.3]).
@@ 3.3.2. Explicit Waldspurger formula
The main result of this subsection is the following.
@@
Let $(\pi,\chi)$ be a self-dual pair as in §3.1 with $\chi$ of conductor ${\mathfrak{c}}=\prod_{v\in S_{0}}{\mathfrak{p}}_{v}^{c_{v}}$. Let $f=\Psi(x_{0}\otimes\varphi)$ be an associated test vector as in Definition 3.8. Suppose that $\mathfrak{N}_{S_{0}}|{\mathfrak{c}}_{S_{0}}$. Then
${\rm N}({\mathfrak{c}}_{1})^{-1}\cdot P(\varsigma^{({\mathfrak{c}})},\varphi,\chi)^{2}=$ $\frac{\Gamma_{\Sigma}(k)\langle{f,f}\rangle_{\widehat{R}^{\times}}}{\pi^{k}(\phi,\phi)_{U_{0}(\mathfrak{N})}}\sqrt{|D_{K}|}\cdot L^{(\Sigma_{0})}(\frac{1}{2},\pi_{K}\otimes\chi)\cdot$
@@
    ${\mathfrak{c}}_{1}=\prod_{v\in S_{0}\backslash S^{-}}{\mathfrak{p}}_{v}^{c_{v}}$,
@@
    $\kappa_{{\mathfrak{c}}_{1}}=\#\ker\left(\mathrm{Pic}(O)\rightarrow\mathrm{Pic}(O_{K,{\mathfrak{c}}_{1}})\right)$ [^11] It equals $1$ or $2$.,
@@
The above result is a consequence of a general explicit Waldspurger formula [9, Thm. 1.8] as we now describe. In loc. cit. the absolute value square of a toric period appears, and the following analysis relates it to the square of the toric period.
@@
For $\chi\in{\mathcal{X}}_{S_{0}}$ of conductor $\mathfrak{c}$} such that $\mathfrak{N}_{S_{0}}|{\mathfrak{c}}_{S_{0}}$, we have
$P(J\varsigma^{({\mathfrak{c}})}\tau,\varphi,\chi)=P(\varsigma^{({\mathfrak{c}})},\varphi,\chi)\frac{\varepsilon(\pi)}{\prod_{{\mathfrak{p}}\in S_{0}}\varepsilon(\pi_{\mathfrak{p}})}.$
@@
$\varsigma_{v}^{-1}J\varsigma_{v}\tau\varphi_{v}=w_{\pi_{v}}\varphi_{v}=\varepsilon(\pi_{v})\varphi_{v},$
where $w_{\pi_{v}}:=\begin{pmatrix}&1\\ \varpi_{v}^{{\mathrm{ord}}_{v}(\mathfrak{N})}\end{pmatrix}\in{\mathrm{GL}}_{2}(F_{v})$ is the Atkin–Lehner operator (cf. [49, Thm. 3.2.2]). If $v|S_{1}^{-}\backslash S_{0}$, then we have $\chi_{v}=1$, $\varsigma_{v}=1$ and ${\mathrm{Hom}}_{K_{v}^{\times}}(\pi_{v},{\mathbb{C}})\neq 0$. Let $P_{v}$ be a basis of the latter space. Then $J$ acts on $P_{v}$ by $\varepsilon(\pi_{v})\varepsilon(B_{v})$ (cf. [44, Thm. 4]). Thus
$P_{v}(J\varsigma_{v}\varphi_{v})=\varepsilon(\pi_{v})\varepsilon(B_{v})P_{v}(\varsigma_{v}\varphi_{v}).$
Now consider the case $v={\mathfrak{p}}\in S_{0}$. Then $\varsigma_{\mathfrak{p}}^{({\mathrm{ord}}_{\mathfrak{p}}{\mathfrak{c}}),-1}J\varsigma_{\mathfrak{p}}^{({\mathrm{ord}}_{\mathfrak{p}}{\mathfrak{c}})}$ stabilises $\varphi$ and so
$P(J\varsigma^{({\mathfrak{c}})},\varphi,\chi)=\prod_{v\in S,v\notin S_{0}}\varepsilon(\pi_{v})\varepsilon(B_{v})P(\varsigma^{({\mathfrak{c}})},\varphi,\chi).$
@@
The following is based on [9, Thm. 1.8].
Taking the test vector in ibid. to be $\varsigma^{({\mathfrak{c}})}f$ with $f=\Psi(x_{0}\otimes\varphi)$, we have
$\frac{P_{\chi}^{0}(\varsigma^{({\mathfrak{c}})}f)P^{0}_{\chi^{-1}}(\varsigma^{({\mathfrak{c}})}\overline{f})}{[O_{K,{\mathfrak{c}}_{1}}^{\times}:O^{\times}]^{2}}{\rm N}({\mathfrak{c}}_{1})^{-1}=\frac{\#\kappa_{{\mathfrak{c}}_{1}}^{2}}{2^{c}}\cdot\frac{\Gamma_{\Sigma}(k)\langle{f,\overline{f}}\rangle_{\widehat{R}^{\times}}}{\pi^{k}(\phi,\phi)_{U_{0}(\mathfrak{N})}}\sqrt{|D_{K}|}\cdot L^{(\Sigma_{0})}(1/2,\pi_{K}\otimes\chi).$
Indeed, this follows from [9, Thm. 1.8] since
@@
    $\nu_{{\mathfrak{c}}_{1}}=\#\kappa_{{\mathfrak{c}}_{1}}[O_{K,{\mathfrak{c}}_{1}}^{\times}:O^{\times}]$
    as in [9, Thm. 1.8].
@@
$P_{\chi}^{0}(\varsigma^{({\mathfrak{c}})}f)=P(\varsigma^{({\mathfrak{c}})},\varphi,\chi).$
In the following, we will use the notation $P_{\chi}^{0}(\varsigma^{({\mathfrak{c}})}f)$ instead of $P(\varsigma^{({\mathfrak{c}})},\varphi,\chi)$ to emphasis the dependence on $x_{0}$. Then in view of the $S$-version of Waldspurger formula [9, Thm. 1.9] we have
@@
$\frac{P_{\chi}^{0}(\varsigma^{({\mathfrak{c}})}f)P_{\chi^{-1}}^{0}(\varsigma^{({\mathfrak{c}})}\tau\overline{f})}{\langle{f,\overline{f}}\rangle}=$ $\frac{P_{\chi}^{0}(\varsigma^{({\mathfrak{c}})}\Psi(x_{0}\otimes\varphi))P_{\chi^{-1}}^{0}(\varsigma^{({\mathfrak{c}})}\tau{\Psi(x_{0}\otimes\varphi)})}{\langle{\Psi(x_{0}\otimes\varphi),\Psi(x_{0}\otimes\varphi)}\rangle_{R}}$
$=$ $\frac{P_{\chi}^{0}(\varsigma^{({\mathfrak{c}})}\Psi(x_{0}\otimes\varphi))P_{\chi}^{0}(J\varsigma^{({\mathfrak{c}})}\tau\Psi(v_{0}\otimes\varphi))}{\langle{\Psi(x_{0}\otimes\varphi),\Psi(x_{0}\otimes\varphi)}\rangle_{R}}$

### v2

@@ 3.2. Test vector
Let $B$ be a quaternion algebra over a number field $F$ and $G=B^{\times}$. Let $\pi$ be a unitary irreducible cuspidal automorphic representation of $G({\mathbb{A}})$ with central character $\omega_{\pi}$ and let $\chi$ be a Hecke character over $K$. The pair $(\pi,\chi)$ is called self-dual if
$\omega_{\pi}\cdot\chi|_{{\mathbb{A}}^{\times}}=1.$
For a self-dual pair $(\pi,\chi)$, we have $\dim{\mathrm{Hom}}_{K_{{\mathbb{A}}}^{\times}}(\pi,\chi^{-1})\leq 1$ and a criterion for the equality is given by (3.2). Following Gross and Prasad [17], $f\in\pi$ is said to be a test vector for ${\mathrm{Hom}}_{K_{{\mathbb{A}}}^{\times}}(\pi,\chi^{-1})$ with respect to an embedding $K_{{\mathbb{A}}}^{\times}\hookrightarrow B_{{\mathbb{A}}}^{\times}$ if any basis of ${\mathrm{Hom}}_{K_{{\mathbb{A}}}^{\times}}(\pi,\chi^{-1})$ takes a non-zero value on $f$. This is essentially a local notion: $f$ is a test vector if it is so locally141414defined analogously for any place. The Waldspurger formula [56, 59] links the Rankin–Selberg $L$-value
@@
with $K^{\times}$-toric period of a test vector in $\pi$ against $\chi$ with respect to an embedding $K\hookrightarrow B$.
Let the setting, and in particular $B$ and $\pi$, be as in §3.1.1. This subsection describes an explicit uniform test vector for the pairs $(\pi,\chi)$ as $\chi$ varies in the family ${\mathcal{X}}_{S_{0}}$ of Hecke characters. The next subsection describes an explicit relation between the $L$-values $L(1/2,\pi_{K}\otimes\chi)$ and toric periods of the test vector.
@@ 3.2.1. Definitions
For a finite place $v$, recall that $M_{0}({\mathfrak{p}}_{v}^{n})_{v}$ is the Eichler of discriminant ${\mathfrak{p}}_{v}^{n}$. Put $U_{0}({\mathfrak{p}}_{v}^{n})_{v}=M_{0}({\mathfrak{p}}_{v}^{n})_{v}^{\times}$. Recall that we have fixed an embedding $\iota:K\hookrightarrow B$ and a family of embeddings $\iota_{\varsigma^{({\mathfrak{r}})}}:K_{{\mathbb{A}}}^{\times}\hookrightarrow G({\mathbb{A}})$ in §2.4.1. Let $S_{1}$ denote the set of places dividing the conductor $\mathfrak{N}$ and $S_{0}$ a finite set of places at which $B$ splits.
@@ 3.2.2. Existence
    If $v\notin S_{0}$, then $R_{v}\cap\iota_{\varsigma_{v}}K_{v}=\iota_{\varsigma_{v}}O_{K_{v}}$, and if $v\in S_{0}$ and $\mathfrak{N}_{v}|{\mathfrak{r}}_{v}={\mathfrak{p}}_{v}^{r_{v}}$, then
    $R_{v}\cap\iota_{\varsigma_{v}^{(r_{v})}}O_{K_{v}}=\iota_{\varsigma_{v}^{(r_{v})}}O_{K_{v},{\mathfrak{r}}_{v}}.$
@@
Part i) just follows from [8, Lem. 3.3].
@@
The first part follows from newform theory and the Tunnell–Saito theorem (cf. [8, Prop. 3.7]). For the second part, we only need to consider the case that $v|\mathfrak{N}^{-}$ but $v\notin S_{0}$, which is the content of [8, Prop. 3.8]. ∎
@@
For any $\chi\in{\mathcal{X}}_{S_{0}}$ of conductor $\mathfrak{c}$}, the vector
$\pi(\varsigma^{({\mathfrak{c}})})f\in\pi$
is a test vector for ${\mathrm{Hom}}_{K_{{\mathbb{A}}}^{\times}}(\pi,\chi^{-1})$ with respect to the embedding $\iota$.
@@
For $v\nmid\mathfrak{N}^{-}$ and $v\notin S_{0}$, note that $B_{v}$ splits and $\chi_{v}$ is unramified. It follws that the local newform $f_{v}$ is a test vector for $\chi_{v}$ and so $f_{v}$ is a test vector. For $v$ such that $v|\mathfrak{N}^{-}$ and $v\nmid S_{0}$, one has $\chi_{v}=1$ and the assertion just follows from the definition. Lastly, for $v\in S_{0}$, the assertion follows from the fact that under the condition $\mathfrak{N}_{v}\mid\mathrm{cond}(\chi_{v})$, the local newform is a test vector for $\chi_{v}$. ∎
@@
Note that the above test vector $f$ only depends on the subset of primes $\{{\mathfrak{q}}\ |\ {\mathfrak{q}}\mid\mathfrak{N}^{-}\ \text{and}\ {\mathfrak{q}}\notin S_{0}\}$. So if $S_{0}$ is coprime to $\mathfrak{N}^{-}$, then $f$ does not depend on $S_{0}$.
@@
For any place ${\mathfrak{q}}|D_{K/F}$ of $F$, suppose that $ord_{{\mathfrak{q}}}(\mathfrak{N}_{\mathfrak{q}})>{\mathrm{ord}}_{{\mathfrak{q}}}(D_{K_{\mathfrak{q}}/F_{\mathfrak{q}}})$. Then ${\rm{N}}(R_{v}^{\times})={\rm{N}}(O_{K_{v}}^{\times})$ for any finite place $v$.
@@
We have the following simple yet key observation.
Lemma 3.12.
Let $M$ be a non-archimedean local field of characteristic 0. Let $D$ be a quaternion algebra over $M$ and $E\subset D$ be a quadratic subalgebra over $M$. Let $\pi$ be an irreducible admissible unitary representation of $D^{\times}$ and let $\chi$ be a character of $E^{\times}$ such that $\dim_{{\mathbb{C}}}{\mathrm{Hom}}_{E^{\times}}(\pi,\chi^{-1})=1$. Let $f\in\pi$ be a test vector151515in the sense of [17] for ${\mathrm{Hom}}_{E^{\times}}(\pi,\chi^{-1})$ of level $U$ for some open compact subgroup $U$ of $D^{\times}$. Let $\overline{U}$ denote the image of $U$ in $D^{\times}/M^{\times}$. Then for any open compact subgroup $V$ of $E^{\times}/M^{\times}$,
$h:=\frac{1}{\#(V/V\cap\overline{U})}\sum_{t\in V/V\cap\overline{U}}\chi(t)\pi(t)f\in\pi$
is also a test vector for ${\mathrm{Hom}}_{E^{\times}}(\pi,\chi^{-1})$ and so non-zero. Furthermore, for any basis $P\in{\mathrm{Hom}}_{E^{\times}}(\pi,\chi^{-1})$, we have
$P(h)=P(f).$
Proof.
Let $T_{\chi^{-1}}$ be the maximal quotient of $\pi$ on which $E^{\times}$ acts via $\chi^{-1}$, which is non-zero by assumption. Then each element in ${\mathrm{Hom}}_{E^{\times}}(\pi,\chi^{-1})$ factors through the natural projection map ${\mathrm{pr}}:\pi\rightarrow T_{\chi^{-1}}$. Note that the map $\mathrm{pr}$} is $E^{\times}$-equivalent and $V/V\cap\overline{U}$ is finite, thus ${\mathrm{pr}}(h)={\mathrm{pr}}(f)$ and the result follows.
Remark 3.13.
Our mod $\ell$ non-vanishing of Hecke $L$-values relies on mod $\ell$ analogue of the above result as in Lemma 4.8.
@@ 3.3.1. Setting
Let $B$ be a quaternion algebra over a number field $F$ and $G=B^{\times}$. Let $\pi$ be an irreducible cuspidal automorphic representation of $G({\mathbb{A}})$ with trivial central character and $\sigma$ its base change to ${\mathrm{GL}}_{2}({\mathbb{A}})$. Let $K$ be a quadratic field extension of $F$ with an embedding $\iota:K\hookrightarrow B$ and $\chi$ a finite order Hecke character over $K$ such that
@@
with the Rankin–Selberg $L$-value $L(1/2,\pi_{K}\otimes\chi)$. Here $dt$ denotes the Tamagawa measure with volume $2L(1,\eta_{K/F})$.
@@
For $\phi$ the ${\mathrm{GL}}_{2}(F)$-newform as normalised in [8, Thm. 1.5] of level
@@
associated to $\pi$, the Petersson norm $(\phi,\phi)_{U_{0}(\mathfrak{N})}$ is defined by the inner product on the Hilbert modular variety ${\mathrm{GL}}_{2}(F)\backslash{\mathbb{H}}^{\Sigma}\times{\mathrm{GL}}_{2}(\widehat{F})/(\widehat{F}^{\times}U_{0}(\mathfrak{N}))$ arising from the invariant measure $dxdy/y^{2}$ on each component $\mathbb{H}$}, which denotes the upper half plane.
Let $\iota:K\hookrightarrow B$ be the embedding as in §3.2.1. Recall that $\widehat{R}\cap\iota_{\varsigma^{({\mathfrak{r}})}}\widehat{O}_{K}=\iota_{\varsigma^{({\mathfrak{r}})}}\widehat{O}_{K,{\mathfrak{r}}}$. For a Hecke character $\chi\in{\mathcal{X}}_{S_{0}}$ of conductor $\mathfrak{r}$}, the toric period $P_{\chi}(\varsigma^{({\mathfrak{r}})}f)$ with respect to embedding $\iota$ is essentially given by [^16] Note that it equals $P_{\chi}^{0}(\pi(\varsigma^{({\mathfrak{r}})})f):=\sum_{[a]\in{\rm P}_{\mathfrak{r}}}\chi(a)\langle{x_{0},\pi(\varsigma^{({\mathfrak{r}})})\varphi(\iota(a))}\rangle_{k-2\Sigma}$ in [8, Thm. 1.8].
$P(\varsigma^{({\mathfrak{r}})},\varphi,\chi):=\sum_{[a]\in{\rm P}_{\mathfrak{r}}}\chi(a)\langle{x_{0},\varphi(x_{\mathfrak{r}}(a))}\rangle_{k-2\Sigma},\quad f=\Psi(x_{0}\otimes\varphi)$ (3.4)
where ${\rm P}_{\mathfrak{r}}=K^{\times}\widehat{F}^{\times}\backslash\widehat{K}^{\times}/\widehat{O}_{K,{\mathfrak{r}}}^{\times}$ (cf. [8, Lem. 2.3]).
@@ 3.3.2. Explicit Waldspurger formula
The main result of this subsection is the following Waldspurger formula for self-dual pairs $(\pi,\chi)$ as $\chi$ varies in anticyclotomic family with restricted places of ramification.
@@
Let $(\pi,\chi)$ be a self-dual pair as in §3.1 with $\chi\in{\mathcal{X}}_{S_{0}}$ of conductor ${\mathfrak{r}}=\prod_{v\in S_{0}}{\mathfrak{p}}_{v}^{r_{v}}$. Let $f=\Psi(x_{0}\otimes\varphi)$ be an associated test vector as in Definition 3.8. Then
${\rm N}({\mathfrak{r}}_{1})^{-1}\cdot P(\varsigma^{({\mathfrak{r}})},\varphi,\chi)^{2}=$ $\frac{\Gamma_{\Sigma}(k)\langle{f,f}\rangle_{\widehat{R}^{\times}}}{\pi^{k}(\phi,\phi)_{U_{0}(\mathfrak{N})}}\sqrt{|D_{K}|}\cdot L^{(\Sigma_{0})}(\frac{1}{2},\pi_{K}\otimes\chi)\cdot$
@@
    $L^{(\Sigma_{0})}(\frac{1}{2},\pi_{K}\otimes\chi)=L(\frac{1}{2},\pi_{K}\otimes\chi)\prod_{v\in S_{0}}L(\frac{1}{2},\pi_{K_{v}}\otimes\chi_{v})^{-1}$ is the incomplete L-value,
  • $\bullet$
    ${\mathfrak{r}}_{1}=\prod_{{\begin{subarray}{c}v\in S_{0}\\ \text{non-split}\end{subarray}}}{\mathfrak{p}}_{v}^{r_{v}}$,
@@
    $\kappa_{{\mathfrak{r}}_{1}}=\#\ker\left(\mathrm{Pic}(O)\rightarrow{\rm P}_{{\mathfrak{r}}_{1}}\right)$ [^17] It equals $1$ or $2$.,
@@
The above result is a consequence of a general explicit Waldspurger formula [8, Thm. 1.8]. In loc. cit. the absolute value square of a toric period appears, and the following analysis relates it to the square of the toric period.
@@
For $\chi\in{\mathcal{X}}_{S_{0}}$ of conductor $\mathfrak{r}$} such that $\mathfrak{N}_{S_{0}}|{\mathfrak{r}}_{S_{0}}$, we have
$P(J\varsigma^{({\mathfrak{r}})}\tau,\varphi,\chi)=P(\varsigma^{({\mathfrak{r}})},\varphi,\chi)\frac{\varepsilon(\pi)}{\prod_{{\mathfrak{p}}\in S_{0}}\varepsilon(\pi_{\mathfrak{p}})}.$
@@
$\pi_{v}(\varsigma_{v}^{-1}J\varsigma_{v}\tau)\varphi_{v}=\pi_{v}(w_{\pi_{v}})\varphi_{v}=\varepsilon(\pi_{v})\varphi_{v},$
where $w_{\pi_{v}}:=\begin{pmatrix}&1\\ \varpi_{v}^{{\mathrm{ord}}_{v}(\mathfrak{N})}\end{pmatrix}\in{\mathrm{GL}}_{2}(F_{v})$ is the Atkin–Lehner operator (cf. [47, Thm. 3.2.2]). If $v|S_{1}^{-}\backslash S_{0}$, then we have $\chi_{v}=1$, $\varsigma_{v}=1$ and ${\mathrm{Hom}}_{K_{v}^{\times}}(\pi_{v},{\mathbb{C}})\neq 0$. Let $P_{v}$ be a basis of the latter space. Then $J$ acts on $P_{v}$ by $\varepsilon(\pi_{v})\varepsilon(B_{v})$ (cf. [42, Thm. 4]). Thus
$P_{v}(\pi_{v}(J\varsigma_{v})\varphi_{v})=\varepsilon(\pi_{v})\varepsilon(B_{v})P_{v}(\pi_{v}(\varsigma_{v})\varphi_{v}).$
Now consider the case $v={\mathfrak{p}}\in S_{0}$. Then $\varsigma_{\mathfrak{p}}^{({\mathrm{ord}}_{\mathfrak{p}}{\mathfrak{r}}),-1}J\varsigma_{\mathfrak{p}}^{({\mathrm{ord}}_{\mathfrak{p}}{\mathfrak{r}})}$ stabilises $\varphi$ and so
$P(J\varsigma^{({\mathfrak{r}})},\varphi,\chi)=\prod_{v\in S,v\notin S_{0}}\varepsilon(\pi_{v})\varepsilon(B_{v})P(\varsigma^{({\mathfrak{r}})},\varphi,\chi).$
@@
The following is based on [8, Thm. 1.8].
Taking the test vector in ibid. to be $\pi(\varsigma^{({\mathfrak{r}})})f$ with $f=\Psi(x_{0}\otimes\varphi)$ and letting for $P_{\chi}^{0}(\pi(\varsigma^{({\mathfrak{r}})})f):=\sum_{[a]\in{\rm P}_{\mathfrak{r}}}\chi(a)\langle{x_{0},\pi(\varsigma^{({\mathfrak{r}})})\varphi(\iota(a))}\rangle_{k-2\Sigma}$, we have
$\frac{P_{\chi}^{0}(\pi(\varsigma^{({\mathfrak{r}})})f)P^{0}_{\chi^{-1}}(\overline{\pi}(\varsigma^{({\mathfrak{r}})})\overline{f})}{[O_{K,{\mathfrak{r}}_{1}}^{\times}:O^{\times}]^{2}}{\rm N}({\mathfrak{r}}_{1})^{-1}=\frac{\#\kappa_{{\mathfrak{r}}_{1}}^{2}}{2^{c}}\cdot\frac{\Gamma_{\Sigma}(k)\langle{f,\overline{f}}\rangle_{\widehat{R}^{\times}}}{\pi^{k}(\phi,\phi)_{U_{0}(\mathfrak{N})}}\sqrt{|D_{K}|}\cdot L^{(\Sigma_{0})}(1/2,\pi_{K}\otimes\chi).$
Indeed, this follows from [8, Thm. 1.8] since
@@
    $\nu_{{\mathfrak{r}}_{1}}=\#\kappa_{{\mathfrak{r}}_{1}}[O_{K,{\mathfrak{r}}_{1}}^{\times}:O^{\times}]$
    as in [8, Thm. 1.8].
@@
$P_{\chi}^{0}(\pi(\varsigma^{({\mathfrak{r}})})f)=P(\varsigma^{({\mathfrak{r}})},\varphi,\chi).$
In the following, we will use the notation $P_{\chi}^{0}(\pi(\varsigma^{({\mathfrak{r}})})f)$ instead of $P(\varsigma^{({\mathfrak{r}})},\varphi,\chi)$ to emphasis the dependence on $x_{0}$. Then in view of the $S$-version of Waldspurger formula [8, Thm. 1.9] we have
@@
$\frac{P_{\chi}^{0}(\pi(\varsigma^{({\mathfrak{r}})})f)P_{\chi^{-1}}^{0}(\overline{\pi}(\varsigma^{({\mathfrak{r}})}\tau)\overline{f})}{\langle{f,\overline{f}}\rangle}=$ $\frac{P_{\chi}^{0}(\pi(\varsigma^{({\mathfrak{r}})})\Psi(x_{0}\otimes\varphi))P_{\chi^{-1}}^{0}(\pi(\varsigma^{({\mathfrak{r}})}\tau){\Psi(x_{0}\otimes\varphi)})}{\langle{\Psi(x_{0}\otimes\varphi),\Psi(x_{0}\otimes\varphi)}\rangle_{R}}$
$=$ $\frac{P_{\chi}^{0}(\pi(\varsigma^{({\mathfrak{r}})})\Psi(x_{0}\otimes\varphi))P_{\chi}^{0}(\pi(J\varsigma^{({\mathfrak{r}})}\tau)\Psi(v_{0}\otimes\varphi))}{\langle{\Psi(x_{0}\otimes\varphi),\Psi(x_{0}\otimes\varphi)}\rangle_{R}}$